prettygirl15:

f(x)=2x^2-12x+16 x intercept(s) axis of symmetry

surjithayer:

any idea?

xxDeppressionxx:

surjithayer:

\[f(x)=2x^2-12x+16\] \[=2(x^2-6x+9-9)+16\] \[=2(x-3)^2-18+16\] \[=2(x-3)^2-2\] vertex (3,-2) axis of symmetry x-3=0 or x=3 to find x-intercepts put f(x)=0 \[2x^2-12x+16=0\] or \[x^2-6x+8=0\] \[x^2-2x-4x+8=0\] \[x(x-2)-4(x-2)=0\] \[(x-2)(x-4)=0\] x=2,4 x intercepts are (2,0),(4,0)

prettygirl15:

@surjithayer wrote:
\[f(x)=2x^2-12x+16\] \[=2(x^2-6x+9-9)+16\] \[=2(x-3)^2-18+16\] \[=2(x-3)^2-2\] vertex (3,-2) axis of symmetry x-3=0 or x=3 to find x-intercepts put f(x)=0 \[2x^2-12x+16=0\] or \[x^2-6x+8=0\] \[x^2-2x-4x+8=0\] \[x(x-2)-4(x-2)=0\] \[(x-2)(x-4)=0\] x=2,4 x intercepts are (2,0),(4,0)
thx

surjithayer:

yw